OpenStudy (superhelp101):

How many grams of O2 are needed to react completely with 12.8 grams of K? K + O2 yields K2O Be sure to balance the equation. 2.62 grams 10.4 grams 51.2 grams 62.6 grams

OpenStudy (superhelp101):

i balenced the equation4K + O2 ---> 2K2O

OpenStudy (superhelp101):

@Abmon98

OpenStudy (superhelp101):

@ganeshie8

OpenStudy (superhelp101):

@Luigi0210 @e.mccormick @Eli_Moses @emmigrace222 @JuliusTheGreat @jdoe0001 @JhonyLovesSchool @zepdrix @ZekeWars

OpenStudy (abmon98):

Balance the Chemical Equation 4K+O2-->2K2O Number of Moles=Mass(g)/Molar Mass(g/mol) 12.8/39.0=0.328 moles of K Ratio of Moles of K to O2 4:1 0.328:x moles of oxygen 0.328*1/4=0.08205 moles of oxygen Mass=Molar Mass*Number of Moles Mass=0.08205*(16*2)

OpenStudy (superhelp101):

i got the answer to be the A

OpenStudy (abmon98):

Yes that's right

OpenStudy (superhelp101):

thanks to you!

OpenStudy (abmon98):

Your welcome :')